the distributive oh net income for the Samson estate is 20000 including 3000 of tax-exempt interest income if they executor distribute $8,000 to the beneficiaries in a year of other than the final year what amount have you taken is the distribution deduction
Answer #1

Solution :-

Distribution deduction = Total distribution - (Total distribution * Exempt interest income / Distributable net income)

= 8000 - (8000 * 3000 / 20000)

= 8000 - (8000 * 0.15)

= 8000 - 1200

= $ 6800.

Conclusion :- Distribution deduction in the given question

= $ 6800.